OpenStudy (anonymous):

x+6y=2 (1) x=4-6y (2) someone help please

OpenStudy (anonymous):

no solutions

OpenStudy (anonymous):

how do you determine that

OpenStudy (turingtest):

solve as a system, right?

OpenStudy (turingtest):

sub for x by putting (2) into (1) (4-6y)+6y=2 4=2 which, because that is impossible, means the system has no solution at all (the lines of their graphs don't intersect). So the answer is "NO SOLUTION"

OpenStudy (anonymous):

each equation represents a line the lines are parallel, dont intersect --> no solution (i.e. no pair of x,y values ) that satisfies both simultanously http://www.wolframalpha.com/input/?i=solve%28x%2B6y%3D2+%3B+x%3D4-6y%29

OpenStudy (anonymous):

(4-6y)+6y=2.... 4=2... so no solution..

OpenStudy (anonymous):

rewrite the equations in the form of ax+by+c=0 x+6y-2=0 (1) x+6y-4=0 (2) if \[\frac{a1}{a2}=\frac{b1}{b2}=\frac{c1}{c2}(infinte solutions)\] if \[\frac{a1}{a2}\neq \frac{b1}{b2}(unique solutions)\] if \[\frac{a1}{a2}=\frac{b1}{b2}\neq\frac{c1}{c2}(no solutions)\] in this case \[\frac{a1}{a2}=\frac{1}{1},\frac{b1}{b2}=\frac{\cancel6}{\cancel6}=\frac{1}{1},\frac{c1}{c2}=\frac{-2}{-4}=\frac{1}{2}\] \[\frac{1}{1}=\frac{1}{1}\neq \frac{1}{2}\] \[\therefore,no solutions\]

OpenStudy (turingtest):

Now that I didn't know :)